A sum of $3,500 is invested in two parts. One part brings a return of 5% and the other a return of 8%. The total annual return is $250. Find the amount invested at each rate.
Answer by oscargut(2103)

A sum of $3,500 is invested in two parts. One part brings a return of 5% and the other a return of 8%. The total annual return is $250. Find the amount invested at each rate.
Let C1 the amount invested at 5% and C2 the amount invested at 8% then
C1 + C2 =3500 and C1(0,05)+C2(0,08)=250 then C2=(3500-C1) then
C1(0,05)+(3500-C1)(0,08)=250 then -0,03C1 +280 =250
then -0,03C1=-30
then C1=1000 and C2=2500